3. Determine the general solution of each differential equation. (a) y – 10y + 25y = 0) (b) 2y – 4y +9 = 0) (C) x2y + 3x
Answer #1

3. (a) y -104 +25 y=0 Aunilary equation m² - lom +25=0 m²- 5m-5m +25=0 (m5)² = 0 m=5,5 Y(x) = c, ar + G n ex Y(u) = (C, + Gwe get Pinlo.al se dy 2 DCD-1) dne where D=d dx hen equation reduce to D (D-1) Z + 3 Dz & 4z = 0 (DCD-1) + 3 D+4) z co wherethumbs up please